OpenStudy (natsent06):

________ Simplify: ∛-216x³

OpenStudy (anonymous):

\[\sqrt[3]{-216x^3}=\sqrt[3]{-216}*\sqrt[3]{x^{3}}=-6x\]

OpenStudy (natsent06):

thanks